# Collapse or Thrive? Perils and Promises of Synthetic Data in a Self-Generating World

### v1

@@ 2.2 Model Collapse in Kernel Density Estimation
We next turn to the second generative modeling setting for studying model collapsed introduced by Shumailov et al. (2024): kernel density estimation (KDE). Similar to multivariate Gaussian modeling, we begin with $n$ real data points drawn from an initial probability distribution $p^{(0)}$: $X_{1}^{(0)},...,X_{n}^{(0)}\sim_{i.i.d.}p^{(0)}$. We then iteratively fit KDEs to the data and sample new synthetic data from these estimators. In the Replace setting, we fit the KDE to $n$ data samples from the most recently fit model, whereas in the Accumulate setting, we fit the KDE to all data points from all previous iterations, with the number of points growing linearly as $n(t+1)$:
$\hat{p}_{\text{Replace}}^{(t+1)}(x)$ =$\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\frac{1}{nh}\sum_{j=1}^{n}K\Big{(}\frac{x-X_{j}^{(t)}}{h}\Big{)}$ (9)
$\hat{p}_{\text{Accumulate}}^{(t+1)}(x)$ =$\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\frac{1}{nh(t+1)}\sum_{i=0}^{t}\sum_{j=1}^{n}K\Big{(}\frac{x-X_{j}^{(i)})}{h}\Big{)}$ (10)
@@
where $K$ is the kernel function and $h$ is the bandwidth parameter. We consider a standard Gaussian kernel. For sampling, at each iteration, we draw $n$ new synthetic data points from the fitted kernel density estimators. We evaluate the performance using the negative log-likelihood (NLL) on real held-out test data; lower NLL indicates better performance. For data, we use four standard synthetic datasets from sklearn (Pedregosa et al., 2011): blobs, circles, moons, and swiss roll.
We again observe the same general difference between replacing data and accumulating data (Fig. 2): replacing data causes a rapid increase in NLL as the number of model-fitting iterations increases, indicating that the KDEs are becoming increasingly poor at modeling the true underlying distribution. In contrast, when data accumulate across model-fitting iterations, we observe that the NLL remains relatively stable, suggesting that accumulating data helps maintain the quality of the KDEs.
Despite the apparent empirical similarities to the iterative Gaussian fitting shown in Figure 1, Gaussian KDEs are theoretically distinct. Unlike Gaussian fitting, regardless of whether we accumulate or replace, the test NLL for a Gaussian KDE asymptotically diverges in theory. There are two interesting caveats: (1) when one begins with a small bandwidth, iteratively fitting KDEs can cause the NLL to initially decrease before it diverges due to the effective kernel bandwidth increasing with model-fitting iterations; (2) although accumulating data causes the NLL to diverge asymptotically, this occurs at a rate so glacial that it doesn’t pose a practical concern. If one wishes to prevent the eventual divergence, one can do so by fitting at each iteration with the optimal bandwidth for the number of data, which should be of the form $c(in)^{-1/5}$ in the $i$th model-fitting iteration as long as data accumulates at a constant rate. Practically speaking, one chooses the bandwidth for KDEs based on the number and characteristics of the data, implying that conscientious practitioners should never witness severe model collapse for KDEs in the accumulate case. For details, see Appendix Sec. B.
We also note a surprising discovery: accumulating data can yield NLLs that decrease with additional model-fitting iterations, meaning that training on real and synthetic data yields lower test loss on real data than training on real data alone. While synthetic data has been shown valuable elsewhere, e.g., Jain et al. (2024), we were surprised to discover this behavior in such a simple setting. This behavior is analogous to Mobahi et al. (2020), which demonstrated how self-distillation of linear models can initially improve model performance by acting as increasing regularization in Hilbert space, but if too many iterations take place, the predictor is regularized towards 00 and performance deteriorates.
2.3 Model Collapse in Supervised Finetuning of Language Models
Figure 3: Model Collapse in Supervised Finetuning of Language Models. Finetuning Google’s Gemma2 2B on Nvidia’s HelpSteer 2 dataset demonstrates that model collapse occurs if previous data are replaced after each model-fitting iteration (left), whereas model collapse is avoided if new synthetic data instead accumulate with previous real and synthetic data (right).
We now turn to the third setting for studying model collapse introduced by Shumailov et al. (2024): supervised finetuning of language models. We begin with an instruction following dataset – Nvidia’s HelpSteer2 (Wang et al., 2024) – and finetune a language model before sampling new text data from it. We choose Google’s Gemma2 2B model (Team et al., 2024) because it is high performing and relatively small. For Replace, we fine-tune the $n$-th language model only on data generated by the $(n-1)$ language model. For Accumulate, we instead fine-tune the $n$-th language model on the starting real data plus all the synthetic data sampled from all previous models; thus, the amount of data for Replace is constant $\sim 12.5k$, whereas the amount of data for Accumulate grows linearly $\sim 12.5k*t$. Consistent with our results and with Gerstgrasser et al. (2024), we find that deleting data after each iteration leads to collapse whereas accumulating data avoids collapse (Fig. 3).
@@ 2.3 Model Collapse in Supervised Finetuning of Language Models
Refer to caption Finetuning Language Models on Instruction-Following Data
@@
Figure 4: Model Collapse Under a Fixed Compute Budget. We compare deleting data after each model-fitting iteration (Replace) and accumulating data after each iteration (Accumulate) with a new fixed-compute data paradigm Accumulate-Subsample. In Accumulate-Subsample, real and synthetic data accumulate but are then subsampled so that each model is trained on a constant number of data. Accumulate-Subsample’s test loss on real data deteriorates more quickly than Accumulate’s loss but more slowly than Replace’s loss, and frequently converges, albeit to a higher plateau than Accumulate. These results hold across five settings: multivariate Gaussian modeling, language model instruction finetuning, kernel density estimation, linear regression and language model pretraining.
3 Model Collapse Under a Fixed Compute Budget
Thus far, we have focused on two data paradigms: Replace and Accumulate. As discussed in Sec. 2, Replace is unlikely to be an faithful model of reality because we do not delete the internet after pretraining each model, but one might also argue that Accumulate is similarly unfaithful because Accumulate requires that every new model is trained on (linearly) more data and thus requires more compute than its predecessor. Whether this criticism is valid in practice is unclear, since newer models are trained on increasing data (e.g., 1.4T tokens for Llama 1, 2T tokens for Llama 2, 15T tokens from Llama 3) and increasing GPUs (e.g., 2k GPUs for Llama1, 4K for Llama2, 16k for Llama3 (Goyal, 2024)). Nevertheless, for the sake of understanding the space of possible outcomes and predicting likely outcomes for future generative models, we ask and answer:
Does model collapse occur when data accumulate but models are trained under a fixed compute budget?
@@ 3 Model Collapse Under a Fixed Compute Budget
We call this data paradigm Accumulate-Subsample because data accumulate but are then subsampled to ensure constant data and thus constant compute at each model-fitting iteration. To study whether model collapse occurs in Accumulate-Subsample, we use the same three generative modeling settings we’ve studied (multivariate Gaussian modeling, supervised finetuning of language models and kernel density estimation) plus two new generative modeling settings studied by prior work (Mobahi et al., 2020; Dohmatob et al., 2024a; Gerstgrasser et al., 2024): linear regression and pretraining language models on a GPT3.5/GPT4-generated dataset of kindergarten-level text (Eldan & Li, 2023).
To explain how linear regression can be used as a generative model, we briefly here and direct the reader to prior work (Mobahi et al., 2020; Dohmatob et al., 2024a; Gerstgrasser et al., 2024) for a more thorough description. We begin with our real covariates $X\in\mathbb{R}^{n\times d}$ and true linear relationship $w^{(0)}$. Initializing $\hat{w}^{(0)}=w^{(0)}$, we sample the regression targets as:
$y^{(t)}\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}X\hat{w}^{(t)}+E^{(t)}\quad;\quad E^{(t)}\sim\mathcal{N}(0,\sigma^{2}I_{d})$ (11)
Assuming $X^{T}X$ is full rank, e.g., $n\gg d$, we fit the next linear model using ordinary least squares:
$\hat{w}^{(t+1)}\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}(X^{T}X)^{-1}X^{T}y^{(t)}$ (12)
Following Gerstgrasser et al. (2024), we additionally pretrain sequences of small variants of common large language models – GPT (Radford et al., 2019; Brown et al., 2020) and Llama (Touvron et al., 2023a; b) – on TinyStories (Eldan & Li, 2023), a synthetic dataset of simple short stories; this combination of models, parameters and data was chosen to faithfully study model collapse in as realistic a setting as possible, subject to our limited computational budget.
@@
Across all five generative modeling settings, we find that Accumulate-Subsample’s test loss on real data lies between the test losses of Replace and Accumulate (Fig. 4 Center). Specifically, Accumulate-Subsample (Fig. 4 center) exhibits higher test loss than Accumulate (Fig. 4, Right) but lower test loss than Replace (Fig. 4 Left), showing that the fixed compute budget imposes some cost. In a qualitative difference, test losses on real data typically plateaus for both Accumulate-Subsample and Accumulate, whereas test losses for Replace typically diverge in an apparently unbounded manner. These results collectively tell a consistent story: under more realistic conditions, where data accumulate and compute is bounded, model performance on real test data is unlikely to diverge.
@@ 4 Cardinality of Real Data vs Proportion of Real Data in Mitigating Model Collapse
Figure 5: The Value of Synthetic Data in Supervised Finetuning of Language Models. Finetuning Google’s Gemma 2 2B on Nvidia’s HelpSteer 2 dataset on different combinations of real and synthetic data demonstrates that loss grows with the number of synthetic data. Our results suggest that the test loss depends on both the proportion (p-value$=3.84\times 10^{-16}$) and the cardinality (p-value$=3.54\times 10^{-8}$) of real data. We plot the test loss against the number of real datapoints in the training set (top left). The hue represents the number of synthetic datapoints. Additionally, we display a heatmap demonstrating the effect of the number of real and synthetic datapoints on test loss (top right). We provide a graph of the test loss versus the fraction of real data, where the hue represents the cardinality of the real data (bottom left). Finally, we plot the test loss against the fraction of real datapoints, where the hue represents the number of synthetic datapoints (bottom right).
We conclude by turning to a question asked by Gerstgrasser et al. (2024) that, to the best of our knowledge, remains open:
Which matters more for avoiding model collapse: the cardinality of real data or the proportion of real data? Relatedly, how does the value of synthetic data for reducing test loss on real data depend on the amount of real data?
These questions are highly pertinent to researchers sampling from web-scale data in order to pretrain or finetune language models. We conduct our investigation of this question as follows: First, we perform SFT on the HelpSteer2 dataset for Google’s Gemma 2 2B model. We sample 100k completions from the finetuned model and filter for those that are fewer than 512 tokens in length. This leaves us with over 55,000 remaining completions. We aggregate datasets containing various numbers of real and synthetic synthetic data, which are given in Figure 5, and perform SFT on these datasets starting from the original Gemma 2B model. We record and display the final test loss from this process.
@@
to best capture the relationship between the fraction of real data and the log likelihood. We measured $R^{2}$ values of $0.59$ for the transformed number of real data and $0.34$ for the proportion of real data. We then computed $F$-statistics for the one-term versus two term models involving each of these covariates, which gave us $p$-values of $6.9\times 10^{-25}$ and $4.6\times 10^{-25}$. These statistics suggest that both the proportion and the cardinality of real data have a statistically significant effect on the test loss, and explain a sizable fraction of the variance in the test loss.
Second, we find a difference in the effect that synthetic data has on test loss in high versus low real data regimes. In our experiments specifically, when the number of real data is 1024 or lower, we find that there is an small but non-zero amount of synthetic data that improves the test loss when it is included. This suggests that practitioners fine-tuning with insufficient amounts of real data should consider supplementing with synthetic data to improve model quality.
On the other hand, when real data are plentiful, we find that more synthetic data almost always harms final model quality when the number of real data is held constant. In some cases, datasets containing only real data prove to be more valuable than datasets that contain ten times more real data mixed with synthetic data.
Although these results are preliminary, they raise some interesting questions about the role of synthetic data in SFT that deserve further exploration. In some of our experiments, we achieve better results by removing all synthetic data from the training set than by doubling the amount of real training data. More generally, when constructing datasets subject to cost constraints, these results suggest that the value of removing synthetic or low-quality data can sometimes exceed that of collecting greater volumes high-quality data.
@@ 5 Discussion
Our work sought to clarify and unify understanding of model-data feedback loops. We demonstrated in three new generative modeling settings that accumulating data over time avoids model collapse, whereas replacing data over time induces model collapse. We then demonstrated in five generative modeling settings that even when each model is trained on a fixed compute budget with a mixture of real and synthetic data, model performance does deteriorate more, but still tends to plateau. The consistency of these results across different model types and datasets suggests that this distinction is a general phenomenon, and is not specific to any particular model, dataset, or learning algorithm. Lastly, we explored the value of synthetic data for reducing the test loss on real data and found two different regimes: when real data are plentiful, synthetic data is harmful, but when real data are scarce, there exists an optimal amount of synthetic data that are helpful.
@@
In our view, the data paradigm in which synthetic data accumulates from a host of models in conjunction with a constant influx of real-world data is more realistic. Under such dynamics, where new synthetic data are added to existing real and synthetic data, model collapse appears unlikely. Our experiments take a pessimistic viewpoint, in the sense that our experiments pay no attention to the quality of data, whereas in practice, engineers heavily filter data based on various indicators of data quality, e.g., (Brown et al., 2020; Lee et al., 2023; Wettig et al., 2024; Penedo et al., 2024; Li et al., 2024; Sachdeva et al., 2024); for a recent review, see Albalak et al. (2024).
6 Future Directions
@@
An especially interesting future direction is how to combine synthetic data generation with filtering techniques to enable performant and efficient pretraining at scale using synthetic data. As we saw in kernel density estimation (Fig. 2) and in language model pretraining on TinyStories (Fig. 4), training on accumulating real and synthetic data can yield lower loss on real test data than training on real data alone. Identifying under what conditions, and why, this is possible is a tantalizing prospect.
@@
Our results in Section 4 suggest that removing low-quality synthetic data from model training sets can improve test loss more than gathering additional high-quality data. Developing efficient removal techniques for detrimental data would streamline the model fine-tuning process and produce better alignment. Future work that further studies and quantifies the harm done by lower-quality synthetic data across broader settings could greatly improve the efficiency of model training.

### v4

@@ 2.2 Model Collapse in Kernel Density Estimation
Shumailov et al. (2024) introduced a second generative modeling task-setting: kernel density estimation (KDE). We begin with $n$ real data points drawn from an initial probability distribution $p^{(0)}$: $X_{1}^{(0)},...,X_{n}^{(0)}\sim_{i.i.d.}p^{(0)}$. We then iteratively fit KDEs to the data and sample new synthetic data from these KDEs. In the replace training-workflow, we fit a KDE to $n$ data from the most recently fit model, whereas in the accumulate training-workflow, we fit a KDE to data from all previous iterations, with the available dataset growing linearly as $n(t+1)$:
$\hat{p}_{\text{Replace}}^{(t+1)}(x)$ =$\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\frac{1}{nh}\sum_{j=1}^{n}K\Big{(}\frac{x-X_{j}^{(t)}}{h}\Big{)}$
$\hat{p}_{\text{Accumulate}}^{(t+1)}(x)$ =$\operatorname{\stackrel{{\scriptstyle\text{def}}}{{\;=\;}}}\frac{1}{nh(t+1)}\sum_{i=0}^{t}\sum_{j=1}^{n}K\Big{(}\frac{x-X_{j}^{(i)})}{h}\Big{)}$
@@
where $K$ is the kernel function and $h$ is its bandwidth parameter. We consider a standard Gaussian kernel. For sampling, at each iteration, we draw $n$ new synthetic data points from the fitted KDEs. We evaluate the performance using the negative log-likelihood (NLL) on real held-out test data; lower NLL indicates better performance. We use four standard synthetic datasets from sklearn (Pedregosa et al., 2011): blobs, circles, moons, and swiss roll.
Our results validate Claims 1 and 2 of Gerstgrasser et al. (2024) in the KDE task-setting. (Fig. 2): replace causes a rapid increase in NLL as iterations increase, indicating that the KDEs are becoming increasingly poor at modeling the real data distribution. In contrast, under accumulate, the NLL of real test data remains relatively stable, demonstrating that this training-workflow avoids model collapse. Surprisingly, accumulating data can yield lower negative log likelihoods on held-out test data that decrease with additional model-fitting iterations. While synthetic data are valuable elsewhere - e.g., Jain et al. (2024), Mobahi et al. (2020) - their value here is intuitive: the synthetic data ”fills in” the gaps between training data points and better approximate the distribution. For the simplest ‘out of the box’ KDE, test losses may diverge (although very slowly) even under accumulate; but, as explained in the Appendices, with sufficient care in setting up the bandwidth selector of the KDE, accumulate test losses will not diverge. By contrast, replace test losses will diverge regardless of the bandwidth selector.
@@
Refer to caption Instruction Finetuning of Language Models
@@
Figure 4: Model Collapse Under a Fixed Compute Budget. We compare deleting data after each model-fitting iteration (replace) and accumulating data after each iteration (accumulate) with a new fixed-compute data paradigm accumulate-subsample. In accumulate-subsample, real and synthetic data accumulate but are then subsampled so that each model is trained on a constant number of data. Accumulate-subsample’s test loss on real data deteriorates more quickly than accumulate’s loss but more slowly than replace’s loss, and frequently converges, albeit to a higher plateau than accumulate. These results hold across five task-settings: multivariate Gaussian modeling, language model instruction finetuning, kernel density estimation, linear regression and language model pretraining.
2.3 Model Collapse in Supervised Fine-tuning of Language Models
We now turn to the third setting for studying model collapse introduced by Shumailov et al. (2024): supervised fine-tuning of language models. Our chosen ‘real data’ is Nvidia’s HelpSteer2 (Wang et al., 2024) instruction-following dataset; we will iteratively fine-tune a language model before sampling new text data from it. We chose Google’s Gemma2 models (Team et al., 2024); they are relatively high performing and relatively small. For replace, we fine-tune the $t$-th language model only on data generated by the $(t-1)$ language model; for accumulate, we instead fine-tune the $t$-th model on the initial real data plus all synthetic data sampled from models $11,\dots,t-$ 1. Thus, the amount of data for replace is constant $\sim 12.5\mbox{\sc k}$, whereas the amount of data for accumulate grows linearly $\sim 12.5\mbox{\sc k}\cdot t$. We again find that deleting data after each iteration leads to collapse whereas accumulating data avoids collapse (Fig. 3).
3 Collapse Under a Fixed Compute Budget?
Thus far, we have focused on two training-workflows: replace and accumulate. As discussed in Sec. 2, replace is unrealistic as a model of internet evolution: we cannot delete and regenerate the internet after pretraining each successive model. Accumulate might also be unrealistic because accumulate trains each successor model on ever-expanding data and increasing compute. For the sake of predicting likely outcomes for future generative models, we ask and answer: Does model collapse occur when data accumulate, while each model training uses a fixed compute budget?
@@ 3 Collapse Under a Fixed Compute Budget?
We call this training-workflow accumulate-subsample since data accumulate but at each model-fitting iteration are subsampled to a fixed dataset size. We study the same three generative modeling task-settings as above, plus two new task-settings from other prior work (Mobahi et al., 2020; Dohmatob et al., 2024a; Gerstgrasser et al., 2024): linear regression and pretraining language models on a GPT3.5/GPT4-generated dataset of kindergarten-level text (Eldan & Li, 2023).
@@
Across all five generative models, accumulate-subsample’s test loss on real data lies between the test losses of Replace and Accumulate (Fig. 4). Specifically, accumulate-subsample exhibits higher test loss than accumulate but lower test loss than replace, showing that the fixed compute budget imposes some performance penalty. In a qualitative difference, test losses on real data typically plateau for both accumulate-subsample and accumulate, while test losses for replace typically diverge in an apparently unbounded manner. Thus, modifying accumulate to the more compute-realistic workflow accumulate-subsample, the threat of model collapse is still avoided.
Figure 5: The Value of Synthetic Data in Supervised Fine-tuning of Language Models. Fine-tuning Google’s Gemma 2 2B on Nvidia’s HelpSteer 2 dataset on different combinations of real and synthetic data. We observe that when the number of real data is small, supplementing with synthetic data can improve test loss. With sufficient real data, adding synthetic data degrades performance.
@@ 4 Cardinality of Real Data vs Proportion of Real Data in Mitigating Model Collapse
We conclude by turning to a key open question: Which matters more for avoiding model collapse: the cardinality of real data or the proportion of real data? Relatedly, how does the value of synthetic data for reducing test loss on real data depend on the amount of real data?
These questions are highly pertinent to researchers sampling from web-scale data in order to pretrain or finetune language models. To explore these questions, we fine-tune Google’s Gemma 2 2B model using the HelpSteer2 dataset. We generate 100k completions from the fine-tuned model and filter out those exceeding 512 tokens, leaving 55,000 usable samples. We then construct datasets with varying mixes of real and synthetic data and fine-tune Gemma 2B on each. The final test loss for each configuration is recorded (Fig. 5).
@@
to best capture the relationship between the fraction of real data and the log likelihood. We measured $R^{2}$ values of $0.59$ for the transformed number of real data and $0.34$ for the proportion of real data. We next computed $F$-statistics for the one-term versus two-term models involving each of these covariates, which gave us $p$-values of $6.9\times 10^{-25}$ and $4.6\times 10^{-25}$, respectively. These statistics suggest that both the proportion and the cardinality of real data have a statistically significant effect on the test loss, and explain a sizable fraction of the variance in the test loss.
Second, we find a difference in the effect of synthetic data on test loss comparing high- versus low- real-data regimes. In our experiments, when the number of real data is 1024 or lower, we find that there is a small but non-zero amount of synthetic data that improves the test loss when it is included. This optimal number of synthetic datapoints appears to be near 1024. This suggests that practitioners fine-tuning with insufficient amounts of real data should consider supplementing with synthetic data to improve model quality. On the other hand, when real data are plentiful (more than 1024 datapoints), we find that more synthetic data almost always degrades final model quality when the number of real data is held constant. When there are more than 1024 real datapoints, datasets containing only real data prove to be more valuable than datasets that contain ten times more real data mixed with synthetic data.
These results raise interesting questions about the role of synthetic data in SFT that merit exploration. We can sometimes achieve better results by removing all synthetic data from the training set than by doubling the amount of real data. When constructing datasets subject to cost constraints, these results suggest that removing synthetic or low-quality data can sometimes bring more value than collecting greater volumes of high-quality data. These experimental results bear comparison with prior theoretical work of Dohmatob et al. (2024b), cf. Corollary 3.3 et seq.
@@ 5 Discussion
We studied the risks of model collapse and identified two pathways to containment. We demonstrated in three new generative modeling settings that accumulating data over time avoids model collapse, whereas replacing data over time induces model collapse. In five generative modeling settings, we demonstrated that using a modified accumulate workflow - where each model trains on a fixed-size sample drawn from all available real and synthetic data - model performance deteriorates, but still tends to plateau. The consistency of results across model types and datasets suggests that this distinction is a general phenomenon, and is not specific to any particular model, dataset, or learning algorithm. Lastly, we explored the value of synthetic data for reducing the test loss on real data and found two different regimes: when real data are plentiful, adding synthetic data can be harmful, but when real data are scarce, there exists an optimal amount of synthetic data that should be added.
@@
To us, the most realistic viewpoint on internet evolution assumes synthetic data accumulating from a host of models alongside continuing influx of real-world data. Combining it with our experimental and analytical work, model collapse seems unlikely. Our experiments take a pessimistic viewpoint, in the sense that our experiments pay no attention to the quality of data, whereas in practice, engineers heavily filter data based on various indicators of data quality, e.g., (Brown et al., 2020; Lee et al., 2023; Wettig et al., 2024; Penedo et al., 2024; Li et al., 2024b; Sachdeva et al., 2024); for a recent review, see Albalak et al. (2024).
@@
A promising future direction might combine synthetic data generation with filtering techniques to enable performant and efficient pretraining at scale. As we saw in kernel density estimation (Fig. 2) and in language model pretraining on TinyStories (Fig. 4), training on real and synthetic data accumulated over several iterations can yield lower loss on real test data than training on real data alone. Identifying under what conditions synthetic data can lower test loss during pretraining would be invaluable to industry practitioners.
@@
Our results in Section 4 suggest that removing low-quality synthetic data from model training sets can improve test loss more than gathering additional high-quality data. Developing efficient identification and removal techniques for detrimental data could streamline the model fine-tuning process and produce better alignment.
Acknowledgements
JK acknowledges support from NSF grant number DGE-1656518. RS acknowledges support from Stanford Data Science and from the OpenAI Superalignment Fast Grant. SK acknowledges support by NSF 2046795 and 2205329, the MacArthur Foundation, Stanford HAI, OpenAI and Google Inc.
Impact Statement
This paper advances the understanding of model collapse, a critical issue in the era of synthetic data proliferation. By systematically analyzing different data evolution paradigms, our findings provide actionable insights into mitigating the risks associated with training generative models on recursively generated data.
The societal implications of our work are substantial. If left unaddressed, model collapse could lead to the degradation of AI-generated content, reducing the reliability of machine learning models deployed across industries, including healthcare, finance, and education. Our results highlight that while synthetic data can be beneficial under controlled accumulation, indiscriminate use may lead to severe model degradation. This has ethical ramifications for AI deployment, as biased or deteriorating models could reinforce misinformation, amplify biases, or reduce transparency in decision-making systems.
Additionally, our study underscores the importance of responsible data curation and compute-aware training strategies. As AI systems become increasingly self-referential, policymakers and practitioners must develop strategies to balance synthetic and real data, ensuring sustainable model performance. Our work informs these discussions by demonstrating that synthetic data, if managed appropriately, can enhance rather than harm AI models.
While our findings primarily contribute to advancing machine learning theory, we encourage continued research into the ethical governance of AI-generated datasets, particularly as reliance on synthetic data expands. Future work should explore mechanisms for filtering low-quality synthetic data and designing robust training paradigms to prevent long-term degradation of AI capabilities.
Appendix A Related Work
The limitations of using AI-generated images to train other image models have been well-documented since 2022 (Hataya et al., 2023). Shumailov et al. (2023) most prominently sounded the alarm about synthetic data for training language models by showing that a model trained repeatedly on its own outputs exhibits severely denigrated quality. Shumailov’s theory and empirical work were quickly extended to many new settings Alemohammad et al. (2024); Bertrand et al. (2024); Dohmatob et al. (2024b; a); Marchi et al. (2024).
Within the model collapse literature, a variety of data dynamics have been studied, which vary in how “real” data are discarded or retained, how “synthetic” data are generated, and how each is (or is not) incorporated into future training sets Martínez et al. (2023); Mobahi et al. (2020); Dohmatob et al. (2024a). A common feature of many of these is that at least some real data are discarded, often because total dataset size is kept constant across model-fitting iterations. However, Gerstgrasser et al. (2024) note that this may not represent real-world dynamics, and that model collapse is avoided when data accumulate. What is not clear, however, is whether this claim holds universally, including in the specific settings studied in other prior work. We help close this gap by extending Gerstgrasser’s empirical and theoretical analyses to several of these settings.
Where model collapse can be seen as studying a worst-case scenario, it has also been observed that some kinds of synthetic data have a positive effect. Dohmatob et al. (2024b) and Jain et al. (2024) find that certain amounts of synthetic data can improve model performance, and Ferbach et al. (2024b) suggest that with curation, self-consuming loops can improve alignment with human preferences. A growing literature on how to filter and harness synthetic data has achieved impressive results on a variety of benchmarks Zelikman et al. (2024); Li et al. (2024a); Yang et al. (2024), raising interesting questions about the limits of when unfiltered synthetic data can help. In this vein, we answer a question posed by Gerstgrasser et al. (2024): does the proportion or the raw amount of real data in a mixed training set have a greater impact on test loss? In the process, we find that proportionally small amounts of synthetic data can improve test loss when real data is scarce.